OpenStudy (anonymous):

Help please Solve for x: −3|x + 7| = −12

pooja195 (pooja195):

Divide both sides by -3. −3(|x+7|)/−3=−12/−3

pooja195 (pooja195):

|x+7|=4 now solve

pooja195 (pooja195):

x+7=-4 x+7=4 solve for x

OpenStudy (anonymous):

thank you

pooja195 (pooja195):

:)

OpenStudy (anonymous):

And @pooja195 saves the day again :). Also, Welcome to Open Study @Usmc2017 have fun.